OpenStudy (anonymous):

The sales price of a car is $12,590.00, which is 20 percent off the original price. What is the original price? $14,990.90 $14,310.40 $15,737.50 $15,290.70

OpenStudy (anonymous):

12590.00 + .20n = n

OpenStudy (jdoe0001):

if 20% has been discounted and came down to $12,590.00 that means that $12,590.00 is only the 80% of the original price the original price will be the 100% thus \(\large \begin{array}{ccllll} price&\% \\\hline\\ 12,590.00&80\\ x&100 \end{array}\implies \cfrac{12,590.00}{x}=\cfrac{80}{100}\)

OpenStudy (anonymous):

can you just solve it please?

OpenStudy (anonymous):

Jdoe0001

OpenStudy (anonymous):

Sorry, my original equation was wrong. It's supposed to be 12590.00 -.2n = n

OpenStudy (jdoe0001):

well... I assume that's what you're meant to do.....what would that give you?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

anyhow just cross-multiply \(\bf \cfrac{12,590.00}{{\color{brown}{ x}}}=\cfrac{80}{{\color{brown}{ 100}}}\implies {\color{brown}{ 100}}\cdot 12,590.00=80\cdot {\color{brown}{ x}}\)

OpenStudy (anonymous):

i cant get the correct answer?

OpenStudy (anonymous):

$15,737.50

OpenStudy (anonymous):

if you dont think its correct put it in a calculator and subtract 20%

OpenStudy (mitu12):

You don't know the original price, so think of that as x. You know that 80% of that number (because 100% - 20% = 80%) is 12,590. In math, "of" means multiply, and "is" means equal. So, we have this equation: .8x = 12,590. To solve for x, divide both sides by .8, and you get the answer: $15,737.50.

OpenStudy (anonymous):

thank uuuu mitu 12

OpenStudy (mitu12):

no problem

OpenStudy (anonymous):

Two angles of a triangle each measure 70 degrees. What is the measure of the third angle in degrees? 80 degrees 100 degrees 120 degrees 40 degrees

OpenStudy (anonymous):

mitu 12 ?